# Globally Rigid Convex Braced Polygons

### v1

@@ Abstract
Determining the global rigidity of a bar framework is known to be a difficult problem. Here we propose a class of frameworks in the plane, braced polygons, that may be globally rigid and are analogous to convex polytopes in $3$-space that are rigid by Cauchy’s Theorem in 1813. We show that all the strictly convex configurations of a given graph, corresponding to a braced polygon (see Section 2), are infinitesimally rigid if and only if they are all globally rigid in the plane in the class of convex braced polygons, Theorem 2.1. Furthermore, we provide techniques that, for many graphs, prove that all strictly convex configurations of the graph of a braced polygon are globally rigid in the plane, and other techniques to find instances when there is a strictly convex braced polygon that is not globally rigid in the plane. As a test case, we use those techniques to determine when a braced polygon is globally rigid or not, when there are $7$ vertices, with $12$ edges in total, and we know that the underlying graph is globally rigid at a generic configuration. Figures 12,14,15. When there are more than $7$ vertices, it looks like global rigidity starts to become difficult again in many cases, but our techniques still work in many cases.
@@ 1.1 History
One of the most fundamental breakthroughs in understanding the rigidity of structures is Cauchy’s 1813 result that a convex triangulated surface is “rigid”. (See [Gruber and Wills(1993)], vol. A 223–271.) This is despite a mistake in his proof of what is now called his “arm lemma”, and this is despite a lack of interest in the subject of rigidity for many years. Nevertheless, the ideas in his proof are correct, and with the passage of time, interest in the subject of rigidity has increased. (See “Rigidity through a projective lens” in [Nixon et al.(2021)Nixon, Schulze, and Whiteley].) The statement of Cauchy’s theorem is clear. There is only one way to assemble a strictly convex triangulated polyhedron, up to congruent copies, that is convex and where vertices are connected with the given correspondence. As a consequence, there is no continuous motion of the polyhedron, where each (triangular) face remains congruent during the motion. The polyhedron is (locally) rigid.
@@
In 1916 Max Dehn [Dehn(1916)] proved that the same strictly triangulated convex polyhedron is infinitesimally rigid in $\mathbb{R}^{3}$. This means that, when there are vectors ${\bf p}^{\prime}=(p^{\prime}_{1},\dots,p^{\prime}_{n})$ attached to the corresponding vertices of the configuration $(G,{\bf p})$ that act as the first derivative of a finite motion, preserving edge (bar) lengths, then they form the derivative of a congruence. This is equivalent to saying that the rank of a matrix, called the rigidity matrix $R({\bf p})$, is maximal, namely $63n-$ 6, where $n$ is the number of vertices of the polyhedron. This also implies that there is no continuous motion, a flex, of the framework that preserves the length of each edge. In the case of a triangulated sphere, each (triangular) face remains congruent during the motion. A priori, Dehn’s theorem looks to be somewhat weaker than Cauchy’s theorem, and, indeed, in 1975 Herman Gluck [Gluck(1975)] showed that there is a relatively easy way to prove Dehn’s infinitesimal theorem using the combinatorial idea of Cauchy’s theorem. It turns out, anyway, that it is easy to show that Cauchy’s theorem directly implies Dehn’s theorem using Proposition 2.2, which is related to some averaging techniques going back to [Connelly and Whiteley(2010)], [Saliola and Whiteley(2004)], and feels like something Minkowski would have done.
@@
Meanwhile, the question of determining the “rigidity” of a bar framework, even in the plane, was considered as a computational question. In this non-necessarily convex setting instead, it was assumed that the coordinates of the configuration ${\bf p}$ are generic, that is, they do not satisfy any non-zero polynomial equation with integer coordinates. So then the (local) rigidity of the framework $(G,{\bf p})$ only depends on the graph $G$ and not on the particular configuration in $\mathbb{R}^{d}$. It was then shown (Geiringer, Laman)[Pollaczek-Geiringer(1926), Laman(1970)] that there is an algorithm to determine generic rigidity for the graph $G$ in $\mathbb{R}^{2}$, and Lovász and Yemini [Lovász and Yemini(1982)], using matriod methods showed that there is a polynomial-time algorithm to compute the generic rigidity of a graph purely combinatorially. Indeed, Bruce Hendrickson and Jacobs [Jacobs and Hendrickson(1997)] have an elementary algorithm, the pebble game, that is easily implemented to compute the generic rigidity of a graph $G$ in the plane.
@@
Another, closely related notion of local rigidity, and infinitesimal rigidity, is the global rigidity of a bar framework $(G,{\bf p})$ in Euclidean space $\mathbb{R}^{d}$. This means that $(G,{\bf p})$ is the only configuration in $\mathbb{R}^{d}$ with the corresponding bars the same length, up to congruent copies. Global rigidity can be much harder to calculate, even when the configuration appears to be generic. But if we are only interested in determining the global rigidity of a framework $(G,{\bf p})$ at generic configurations, things become easier.
Still there is another problem, even if the configuration ${\bf p}$ is known exactly. A test for generic global rigidity (GGR) is that there is an equilibrium stress $\omega$ for a given framework $(G,{\bf p})$ with a corresponding stress matrix $\Omega$ that has maximal rank. The top two frameworks of Figure 1 both have a one-dimensional stress with a stress matrix $\Omega$ that has maximal rank. In this case the rank is $n-d-1=8-2-1=5$, where $n=8$ is the number of vertices, and $d=2$ is the dimension of the ambient space. The frameworks are both infinitesimally rigid in the plane, and the vertices form strictly convex polygons. The two configurations have corresponding edge lengths the same and are not congruent, even though they testify to the generic rigidity of the corresponding graph. When one knows that the graph $G$ is generically globally rigid, in practice, it can be extremely tedious, and at least difficult, to determine the global rigidity of a given framework $(G,{\bf p})$, even in the plane. (In [Hendrickson(1992)] it is shown that this is an NP-complete problem to determine.) It is possible, with a given graph $G$, to find frameworks that are globally rigid in all higher dimensions, starting in a given lower dimension, universal rigidity, as in [Oba and Tanigawa(2023)], but this may not deal with one’s favorite configuration. One could say that when we say that a graph $G$ is generically globally rigid in a given dimension, we do not know what we are talking about.
@@ 1.2 Our results
If one is given a particular tensegrity, or a class of tensegrities, as described below, how can one determine global rigidity? We propose, here, some tools to do that.
@@
In Section 2, we define the class of strictly “convex braced polygons”, which are an open set of configurations which correspond to a particular class of graphs. This makes it easier to test if a particular configuration in that class has a strictly convex configuration that is not globally rigid, given that one can know when it is infinitesimally rigidity. For general graphs, it might be more helpful to use something like the “pure condition” in [White and Whiteley(1983)] to find the critical configurations, where the rigidity matrix $R({\bf p})$ drops rank, for instance. See also [Nixon et al.(2021)Nixon, Schulze, and Whiteley] for a history of some of the classical methods for computing the (infinitesimal) rigidity of bar frameworks.
In Section 3, we show that if a 3-connected braced polygon circuit (in the sense matroid theory) is minimal with respect to the braces, then it always has a strictly convex configuration, where the stress is positve on the outer polygon and negative on all the braces, Theorem 3.6. This result is equivalent to an earlier result of [Geleji and Jordán(2013)] using what they call the “unique interval property” (Definition A.1). Our proof here makes no mention of that property. The result implies global rigidity by a theorem of Connelly[Connelly and Guest(2022), Theorem 5.14.4].
@@
In Section 4, we give some examples that cannot be resolved by the methods in the previous sections, as well as the techniques we can use to determine their global rigidity as well as an averaging technique to determine, in many cases, their non-global rigidity. One very useful method is to use superposition of configurations that can work very efficiently with both the stress matrix $\Omega$ and the rigidity matrix $R({\bf p})$. These methods can solve all cases of 7, or fewer, vertices. See Table 14. We also provide a table of the $7$ vertex strictly convex braced polygons that are unique in the class of strictly convex braced polygons, Figure 15, as mentioned above, in the spirit of Cauchy’s Theorem. In some cases, one can do calculations simply by seeing how configurations cross thresholds without doing computer calculations. See Figure 21 for an example, and Remark 4.1 for an even simpler example to create globally rigid (and universally rigid) configurations. (This reminds one of some of the “catastrophe” techniques in [Heaton and Timme(2022)], although Catastrophe Theory, itself, does not seem helpful.) An example that we cannot solve is also provided(Figure 20), together with other unsolved natural questions.
@@ 2 Convex Braced Polygons
Suppose that a strictly convex polygon in the Euclidean plane has edges with fixed length and some additional edges, also of fixed length, connecting some of the polygon vertices. We call this a convex braced polygon. We are concerned here when this bar framework is rigid in four senses, infinitesimally rigid, convexly rigid, globally rigid, and universally rigid. We call $(G,{\bf p})$ a bar framework, where $G$ is a finite graph whose vertices correspond to a configuration ${\bf p}=(p_{1},\dots,p_{n})$ of points $p_{i}$, $i=1\dots,n$ in some Euclidean space. The edges of $G$ correspond to fixed length bars, but not necessarily of the same length, connecting the vertices of $(G,{\bf p})$. A convex polygon in the plane is strictly convex if each vertex $p_{i}$ is in a line that intersects the polygon only at $p_{i}$, and for a convex braced polygon we assume that the vertices $(p_{1},\dots,p_{n})$ form a convex polygon with each edge $i,i+1$, modulo $n$, corresponding to a boundary edge of the polygon. When we say $(G,{\bf p})$ and $(G,{\bf q})$ have corresponding bars the same length, we mean that for each $i,j$ a bar of $G$, $\|p_{i}-p_{j}\|=\|q_{i}-q_{j}\|$, and we assume that they are both in clockwise order or both in counter-clockwise order.
@@
For a bar framework $(G,{\bf p})$ in $\mathbb{R}^{d}$, and ${\bf p}^{\prime}=(p^{\prime}_{1},\dots,p^{\prime}_{n})$, regarded as a configuration of vectors $p^{\prime}_{i}$ in $\mathbb{R}^{d}$, we say ${\bf p}^{\prime}$ is an infinitesimal flex of $(G,{\bf p})$, when for each edge $i,j$ of $G$
$({\bf p}_{i}-{\bf p}_{j})\cdot({\bf p}_{i}^{\prime}-{\bf p}^{\prime}_{j})=0$
@@
Definition 2.2.
A braced strictly convex polygon $(G,{\bf p})$ in $\mathbb{R}^{2}$ is convexly rigid if every other strictly convex braced polygon $(G,{\bf q})$, whose corresponding edge lengths are the same, is congruent to $(G,{\bf p})$. (Two frameworks are congruent if $\|p_{i}-p_{j}\|=\|q_{i}-q_{j}\|$, for all $1\leq i<j\leq n$.)
@@
Notice that for the graph $G$ of a braced polygon, the ordering of the vertices is understood as part of the graph $G$. With this in mind we have the following:
@@
For a given braced polygon graph $G$, all strictly convex braced polygon frameworks $(G,{\bf p})$ are infinitesimally rigid, if and only if they are all convexly rigid.
One direction of this theorem follows from the following, which is related to ideas of Minkowski and extends easily to tensegrities by keeping track of the inequalities:
@@
For any non-trivial infinitesimal flex ${\bf p}^{\prime}$ of a bar framework $(G,{\bf p})$, whose affine span of the vertices of ${\bf p}$ is all of $\mathbb{R}^{d}$, then $(G,{\bf p}+{\bf p}^{\prime})$ and $(G,{\bf p}-{\bf p}^{\prime})$ have corresponding bar lengths the same, but are not congruent.
Proof. If $(G,{\bf p}^{\prime})$ is an infinitesimal flex of $(G,{\bf p})$, and $i,j$ corresponds to a bar of $G$, then $(p_{i}-p_{j})\cdot(p^{\prime}_{i}-p^{\prime}_{j})=0$, and so
@@
and thus $(G,{\bf p}+{\bf p}^{\prime})$ and $(G,{\bf p}-{\bf p}^{\prime})$ have corresponding bar lengths the same. When ${\bf p}^{\prime}$ is non-trivial, and the affine span of ${\bf p}$ is all of $\mathbb{R}^{d}$, then for some pair of vertices corresponding to $i,j$, $(p_{i}-p_{j})\cdot(p^{\prime}_{i}-p^{\prime}_{j})\neq 0$, and the above calculation shows that $(G,{\bf p}+{\bf p}^{\prime})$ and $(G,{\bf p}-{\bf p}^{\prime})$ are not congruent, since the middle equality, above, is violated. ∎
One can see how this works from Figure 1. Notice that, due to the symmetry involved, the two non-congruent braced polygons on the top are mirror images of each other, but that congruence does not preserve the vertex labeling. This symmetry can be broken with different examples.
Figure 1: The two braced top polygons have corresponding bar lengths the same, but are not congruent. The bottom braced polygon, for the same graph, has an infinitesimal flex, where each non-zero $p^{\prime}_{i}$ is indicated by a small arrow. This provides the two graphs on top.
Next we need some lemmas related to convex polygons.

### v2

@@ Abstract
Determining the global rigidity of a bar framework is known to be a difficult problem. Here we propose a class of frameworks in the plane, convex braced polygons, that may be globally rigid and are analogous to convex polyhedra in $3$-space, which are globally rigid in the class of convex polyhedra by a celebrated theorem of Cauchy from 1813. We show that all the strictly convex plane realisations of the underlying graph of a convex braced polygon are infinitesimally rigid if and only if they are all globally rigid in the class of convex braced polygons. Furthermore, we provide techniques that, for many such graphs, prove that all their strictly convex plane realisations are globally rigid, and other techniques to find instances when there is a strictly convex plane realisation that is not globally rigid.
@@ 1.1 History
One of the most fundamental breakthroughs in understanding the rigidity of structures is Cauchy’s 1813 result that a convex triangulated surface is “rigid”. (See [Gruber and Wills(1993)], vol. A 223–271.) This is despite a mistake in his proof of what is now called his “arm lemma”, and also a lack of interest in the subject of rigidity for many years. Nevertheless, the ideas in his proof are correct, and with the passage of time, interest in the subject of rigidity has increased. (See “Rigidity through a projective lens” in [Nixon et al.(2021)Nixon, Schulze, and Whiteley].) The statement of Cauchy’s theorem is clear. There is only one way to assemble a strictly convex triangulated polyhedron, up to congruent copies, that is convex and where vertex adjacencies are preserved. As a consequence, there is no continuous motion of the polyhedron, where each (triangular) face remains congruent during the motion. The polyhedron is (locally) rigid.
@@
In 1916 Max Dehn [Dehn(1916)] proved that the same strictly convex triangulated polyhedron is infinitesimally rigid in $\mathbb{R}^{3}$. This means that, when there are vectors ${\bf p}^{\prime}=(p^{\prime}_{1},\dots,p^{\prime}_{n})$ attached to the corresponding vertices of the configuration $(G,{\bf p})$ that act as the first derivative of a finite motion, preserving edge (bar) lengths, then they form the derivative of a congruence. This is equivalent to saying that the rank of a matrix, called the rigidity matrix $R({\bf p})$, is maximal, namely $63n-$ 6, where $n$ is the number of vertices of the polyhedron. This also implies that there is no continuous motion, a flex, of the framework that preserves the length of each edge. In the case of a triangulated polyhedron, each (triangular) face remains congruent during the motion. A priori, Dehn’s theorem looks to be somewhat weaker than Cauchy’s theorem, and, indeed, in 1975 Herman Gluck [Gluck(1975)] showed that there is a relatively easy way to prove Dehn’s infinitesimal theorem using the combinatorial idea of Cauchy’s theorem. It turns out, anyway, that it is easy to show that Cauchy’s theorem directly implies Dehn’s theorem using Proposition 2.2 below, which is related to some averaging techniques going back to [Connelly and Whiteley(2010)], [Saliola and Whiteley(2004)], and feels like something Minkowski would have done.
@@
Meanwhile, the question of determining the “rigidity” of a bar framework, even in the plane, was considered as a computational question. In this not-necessarily convex setting, it is often assumed that the coordinates of the configuration ${\bf p}$ are generic, that is, they do not satisfy any non-zero polynomial equation with integer coordinates. In this case, the (local) rigidity of the framework $(G,{\bf p})$ only depends on the graph $G$ and not on the particular configuration in $\mathbb{R}^{d}$. (Geiringer, Laman)[Pollaczek-Geiringer(1926), Laman(1970)] characterised graphs which are generically rigid in $\mathbb{R}^{2}$, and Lovász and Yemini [Lovász and Yemini(1982)], used matriod methods to show there is a combinatorial polynomial-time algorithm to determine the generic rigidity of a graph. Indeed, Bruce Hendrickson and Jacobs [Jacobs and Hendrickson(1997)] have an elementary algorithm, the pebble game, that is easily implemented to compute the generic rigidity of a graph $G$ in the plane.
@@
Another, closely related notion of local rigidity, and infinitesimal rigidity, is the global rigidity of a bar framework $(G,{\bf p})$ in Euclidean space $\mathbb{R}^{d}$. This means that $(G,{\bf p})$ is the only configuration in $\mathbb{R}^{d}$ with the corresponding bars the same length, up to congruent copies. Global rigidity can be much harder to test, but if we are only interested in determining the global rigidity of a framework $(G,{\bf p})$ at generic configurations, the problem becomes more tractable.
Still there is another problem, even if the configuration ${\bf p}$ is known exactly. A test for generic global rigidity (GGR) is that there is an equilibrium stress $\omega$ for a given framework $(G,{\bf p})$ with a corresponding stress matrix $\Omega$ that has maximal rank. The top two frameworks of Figure 1 both have a one-dimensional stress with a stress matrix $\Omega$ that has maximal rank. In this case the rank is $n-d-1=8-2-1=5$, where $n=8$ is the number of vertices, and $d=2$ is the dimension of the ambient space. The frameworks are both infinitesimally rigid in the plane, and the vertices form strictly convex polygons. The two configurations have corresponding edge lengths the same and are not congruent, even though they testify to the generic rigidity of the corresponding graph. When one knows that the graph $G$ is generically globally rigid, in practice, it can be extremely tedious, and at least difficult, to determine the global rigidity of a given framework $(G,{\bf p})$, even in the plane. (In [Hendrickson(1992)] it is shown that this is an NP-complete problem to determine.) It is possible, with a given graph $G$, to find frameworks that are globally rigid in all higher dimensions, starting in a given lower dimension, universal rigidity, as in [Oba and Tanigawa(2024), Oba and Tanigawa(2023)], but this may not deal with one’s favorite configuration. Knowing that a graph is generically globally rigid in a given dimension does not tell us anything about the global rigidity of a particular realisation.
@@ 1.2 Our results
If one is given a particular framework, or a class of frameworks, as described below, how can one determine global rigidity? We propose, here, some tools to do that.
@@
In Section 2, we consider the class of strictly “convex braced polygons”. We show that all the strictly convex plane realisations of the underlying graph of a convex braced polygon are infinitesimally rigid if and only if they are all globally rigid in the class of convex braced polygons (Theorem 2.1). This makes it easier to test if a particular configuration in that class has a strictly convex configuration that is not globally rigid, given that one can know when it is infinitesimally rigid. For general graphs, it might be more helpful to use something like the “pure condition” in [White and Whiteley(1983)] to find the critical configurations, where the rigidity matrix $R({\bf p})$ drops rank, for instance. See also [Nixon et al.(2021)Nixon, Schulze, and Whiteley] for a history of some of the classical methods for computing the (infinitesimal) rigidity of bar frameworks.
In Section 3, we show that, if a graph of a convex braced polygon is minimally 3-connected with respect to its set of braces, then the property that all its strictly convex plane realisations are globally rigid (the property studied in Section 2) is equivalent to the property that every strictly convex plane realization has a proper stress, i.e., a stress which is positive on the boundary edges and negative on the braces (Theorem 3.6). The latter property implies global rigid by a theorem of Connelly[Connelly and Guest(2022), Theorem 5.14.4]. Theorem 3.6 is closely related to an earlier result of Geleji and Jordán [Geleji and Jordán(2013)] on convex braced polygons whose underlying graph is a 3-connected circuit in the generic 2-dimensional rigidity matroid. Geleji and Jordán [Geleji and Jordán(2013)] characterized when such a graph has the property that all its strictly convex realizations have a proper stress. We shall give a new characterization of this property in Theorem 3.7. Geleji and Jordán [Geleji and Jordán(2013)] focused on a special combinatorial property of the graphs of convex braced polygons, called the “unique interval property” (Definition A.1). Our proof of Theorem 3.7 makes no mention of that property, and uses minimal 3-connectivity instead. We will derive a combinatorial relation between minimal 3-connectivity and the unit interval property in the appendix.
@@
In Section 4, we will give some examples that cannot be resolved by the results in the previous sections. We describe some techniques which can be used to verify their global rigidity as well as an averaging technique which can verify non-global rigidity. One very useful method is to use superposition of configurations that can work very efficiently with both the stress matrix $\Omega$ and the rigidity matrix $R({\bf p})$. These techniques can be used to determine whether any graph of a convex braced polygon with at most 7 vertices is globally rigid at all strictly convex realisations in the plane, see Figure 15. We also provide a table of the graphs of convex braced polygons on at most 7 vertices whose strictly convex realisations are globally rigid when restricted to the class of all strictly convex realisations, see Figure 16. In some cases, one can do calculations simply by seeing how configurations cross thresholds without doing computer calculations. See Figures 20 and 21 for an example. (This reminds one of some of the “catastrophe” techniques in [Heaton and Timme(2022)], although Catastrophe Theory, itself, does not seem helpful.) A graph on eight vertices for which we cannot determine whether all strictly convex realisations are globally rigid is also provided (Figure 18), together with other unsolved natural questions.
@@ 2 Convex Braced Polygons
A $d$-dimensional bar framework is a pair $(G,{\bf p})$ , where $G$ is a finite graph whose vertices correspond to a configuration ${\bf p}=(p_{1},\dots,p_{n})$ of points $p_{i}$, $i=1\dots,n$ in $d$-dimensional Euclidean space. The edges of $G$ correspond to fixed length bars, but not necessarily of the same length, connecting the vertices of $(G,{\bf p})$. We will also refer to $(G,{\bf p})$ as a realisation of $G$ in $\mathbb{R}^{d}$. We will say that $(G,{\bf p})$ and $(G,{\bf q})$ have corresponding bars the same length to mean that, for each edge $ij$ of $G$, $\|p_{i}-p_{j}\|=\|q_{i}-q_{j}\|$. Two $d$-dimensional frameworks $(G,{\bf p})$ and $(G,{\bf q})$ are congruent if $\|p_{i}-p_{j}\|=\|q_{i}-q_{j}\|$, for all $1\leq i<j\leq n$.
@@
For a bar framework $(G,{\bf p})$ in $\mathbb{R}^{d}$, and ${\bf p}^{\prime}=(p^{\prime}_{1},\dots,p^{\prime}_{n})$, regarded as a configuration of vectors $p^{\prime}_{i}$ in $\mathbb{R}^{d}$, we say ${\bf p}^{\prime}$ is an infinitesimal flex of $(G,{\bf p})$ if, for each edge $ij$ of $G$, we have
$({\bf p}_{i}-{\bf p}_{j})\cdot({\bf p}_{i}^{\prime}-{\bf p}^{\prime}_{j})=0.$
@@
A convex polygon in the plane is strictly convex if each vertex is in a line that intersects the polygon only at this vertex. Suppose that a (strictly) convex polygon in the plane has edges with fixed length and some additional edges, also of fixed length, connecting some of the polygon vertices. We call this a (strictly) convex braced polygon, and we regard it as a bar framework in the plane. For a convex braced polygon we assume that the vertices $(p_{1},\dots,p_{n})$ form a convex polygon with each edge $i,i+1$, modulo $n$, corresponding to a boundary edge of the polygon. The underlying graph of a convex braced polygon is called a braced polygon graph. It is a Hamiltonian graph with a distinguished Hamilton cycle.
Motivated by Cauchy’s rigidity theorem, we introduce the following rigidity notion for strictly convex braced polygons.
Definition 2.4.
A braced strictly convex polygon $(G,{\bf p})$ in $\mathbb{R}^{2}$ is convexly rigid if every other strictly convex braced polygon $(G,{\bf q})$, whose corresponding edge lengths are the same, is congruent to $(G,{\bf p})$. In addition, we say that a braced polygon graph is convexly rigid if all its strictly convex configurations are convexly rigid.
We are concerned here with the problem of deciding when a braced polygon graph has the property that all its strictly convex realisations are rigid in four senses, infinitesimally rigid, convexly rigid, globally rigid, and universally rigid. The first result in this paper is the following theorem.
@@
For a given braced polygon graph $G$, all strictly convex braced polygon frameworks $(G,{\bf p})$ are infinitesimally rigid if and only if they are all convexly rigid.
One direction of this theorem follows from the following proposition, which is related to ideas of Minkowski.
@@
Let $(G,{\bf p})$ be a bar framework whose points affinely span $\mathbb{R}^{d}$ and ${\bf p}^{\prime}$ be a non-trivial infinitesimal flex of $(G,{\bf p})$. Then $(G,{\bf p}+{\bf p}^{\prime})$ and $(G,{\bf p}-{\bf p}^{\prime})$ have corresponding bar lengths the same, but they are not congruent.
Proof. Suppose $ij\in E(G)$. Then $(p_{i}-p_{j})\cdot(p^{\prime}_{i}-p^{\prime}_{j})=0$, and so
@@
Hence $(G,{\bf p}+{\bf p}^{\prime})$ and $(G,{\bf p}-{\bf p}^{\prime})$ have corresponding bar lengths the same. When ${\bf p}^{\prime}$ is non-trivial, and the affine span of ${\bf p}$ is all of $\mathbb{R}^{d}$, then for some pair of vertices $i,j\in V(G)$, $(p_{i}-p_{j})\cdot(p^{\prime}_{i}-p^{\prime}_{j})\neq 0$, and the above calculation shows that $(G,{\bf p}+{\bf p}^{\prime})$ and $(G,{\bf p}-{\bf p}^{\prime})$ are not congruent, since the middle equality, above, is violated. ∎
Proposition 2.2 is illustrated in Figure 1. Notice that, due to the symmetry involved, the two non-congruent convex braced polygons on the top are mirror images of each other, but that congruence does not preserve the vertex labeling. The mirror symmetry can be broken with different examples.
Figure 1: The two convex braced polygons on the top have corresponding bar lengths the same, but are not congruent. The bottom convex braced polygon has an infinitesimal flex ${\bf p}^{\prime}$, where each non-zero $p^{\prime}_{i}$ is indicated by a small arrow. This gives rise to the two convex braced polygons on the top.
The converse direction of Proposition 2.2 also holds and it is stated as follows. The proof follows by reversing that of Proposition 2.2.
Proposition 2.3.
Let $(G,{\mathbf{p}})$ and $(G,{\mathbf{q}})$ be two non-congruent frameworks having corresponding bar lengths the same. Then ${\mathbf{p}}-{\mathbf{q}}$ is a non-trivial infinitesimal motion of $(G,\frac{{\mathbf{p}}+{\mathbf{q}}}{2})$.
Now let us move to the proof of Theorem 2.1. Proposition 2.2 implies that, if some braced strictly convex polygon $(G,{\mathbf{p}})$ has a non-trivial infinitesimal flex ${\mathbf{p}}^{\prime}$, then $(G,{\mathbf{p}}+\varepsilon{\mathbf{p}}^{\prime})$ is not convexly rigid for a sufficiently small $\varepsilon>0$. This gives one direction of Theorem 2.1. We will verify the converse direction by using Proposition 2.3, but this step requires an additional idea because $(G,\frac{{\mathbf{p}}+{\mathbf{q}}}{2})$ may not be a strictly convex braced polygon even if $(G,{\mathbf{p}})$ and $(G,{\mathbf{q}})$ are. See Figure 4 for an example. Our goal is to show the existance of an appropriate rotation of $(G,{\mathbf{q}})$ such that $(G,\frac{{\mathbf{p}}+{\mathbf{q}}}{2})$ is a strictly convex braced polygon after rotation.
Suppose that $(G,{\mathbf{p}})$ and $(G,{\mathbf{q}})$ are strictly convex braced polygons having corresponding bar lengths the same, and denote ${\mathbf{p}}=(p_{1},\dots,p_{n})$ and ${\mathbf{q}}=(q_{1},\dots,q_{n})$. We prepare a reference ray to be the horizontal ray to the left direction, and for each $i=1,\dots,n$ let $\theta_{i}$ (resp. $\phi_{i}$) be the counterclockwise angle from the reference ray to $p_{i+1}-p_{i}$ (resp. $q_{i+1}-q_{i}$). By rotating $(G,{\mathbf{p}})$ (resp. $(G,{\mathbf{q}})$), we may assume that $\theta_{1}=\phi_{1}=0$. See Figure 2. We can further extend the definition of $\theta_{i}$ (and similarly $\psi_{i}$) over any $i\in\mathbb{Z}$ by choosing $r,t\in\mathbb{Z}$ such that $i=tn+r$ and $1\leq r\leq n$, and then setting $\theta_{i}:=\theta_{r}+2\pi t$.
Figure 2: Angles in a strictly convex polygon.
By strict convexity, we have $\theta_{i}<\theta_{i+1}<\theta_{i}+\pi$ for each $i$, and this inequality is indeed a necessary and sufficient condition for $(G,{\mathbf{p}})$ to be strictly convex.
Consider $(G,\frac{{\mathbf{p}}+{\mathbf{q}}}{2})$ and define the corresponding angles $\psi_{i}\ (i\in\mathbb{Z})$ for $(G,\frac{{\mathbf{p}}+{\mathbf{q}}}{2})$ by the same definition. Then $\psi_{i}=\frac{\theta_{i}+\phi_{i}}{2}$ if $|\theta_{i}-\phi_{i}|$ is less than $\pi$. (If $|\theta_{i}-\phi_{i}|$ is more than $\pi$, then the direction of $\frac{p_{i+1}+q_{i+1}}{2}-\frac{p_{i}+q_{i}}{2}$ becomes opposite to the angle we are measuring between $p_{i+1}-p_{i}$ and $q_{i+1}-q_{i}$. So $\psi_{i}$ becomes $\frac{\theta_{i}+\phi_{i}}{2}-\pi$. See Figure 3.) This implies that, if we can guarantee
$|\theta_{i}-\phi_{i}|<\pi\quad\text{for all $i$},$ (1)
then, for all $i$, $\psi_{i}=\frac{\theta_{i}+\phi_{i}}{2}$ holds and $\psi_{i}<\psi_{i+1}<\psi_{i}+\pi$ follows from $\theta_{i}<\theta_{i+1}<\theta_{i}+\pi$ and $\phi_{i}<\phi_{i+1}<\phi_{i}+\pi$. Thus, $(G,\frac{{\mathbf{p}}+{\mathbf{q}}}{2})$ will be strictly convex.
Our next lemma shows that we can always achieve (1) by rotating $(G,{\mathbf{q}})$. We say that a function $f:\mathbb{R}\rightarrow\mathbb{R}$ is a monotone step function if it is a piecewise constant, upper-semi-continuous function, and non-decreasing.